OpenStudy (anonymous):

Find the length of an arc of 40° in a circle with an 8 inch radius.

OpenStudy (ajprincess):

length=(40/360)*2*pi*r

OpenStudy (anonymous):

\[l = r \times \theta \]

OpenStudy (unklerhaukus):